For a gas reactor containing 10 I2 molecules and 6 Cl2 molecules, answer the questions that follow.

I2(g)+3Cl2(g)------>2ICL3(s)

Based on the balanced chemical equation and the number of reactant molecules, how many reaction cycles will be completed?

How many ICl3 molecules will be produced?

How many I2 molecules will be leftover?

How many Cl2 molecules will be leftover?
Answer #1

I2(g)+3Cl2(g)------>2ICL3(s)

1 molecule of I2 react with 3 molecules of Cl2

10 molecule of I2 react with = 3*10 /1   =30 molecules of Cl2

Cl2 is limiting reactant

3 molecules of Cl2 react with I2 to gives 2 molecules of ICl3

6 molecules of Cl2 react with I2 to gives = 2*6/3    = 4 molecules of ICl3

3 molecules of Cl2 react with 1 molecules of I2

6 molecules of Cl2 react with = 1*6/3 = 2 molecules of I2 is required.

I2 is excess reactant

The molecules of excess reactant left over = 10-2 = 8 molecules of I2

Cl2 is limiting reactant . It is completely participate in chemical reaction.

The molecules are left over after complete the reaction is zero